Using MVT theorem to prove that

x/(1+x^2) < tan (^-1) x < x
1
Expert's answer
2021-11-01T10:16:41-0400

Prove the inequality for "x>0" (otherwise it doesn’t hold)

"\int\limits_{0}^{x}{\frac{1}{1+{{t}^{2}}}dt}={{\tan }^{-1}}x"

By MVT

"\int\limits_{0}^{x}{\frac{1}{1+{{t}^{2}}}dt}=\frac{1}{1+{{\xi }^{2}}}x,\xi \in \left( 0,x \right)"

We have

"\frac{1}{1+{{\xi }^{2}}}\cdot x<1\cdot x=x \\ 

\frac{1}{1+{{\xi }^{2}}}\cdot x>\frac{x}{1+{{x}^{2}}} \\"

Thus

"\frac{x}{1+{{x}^{2}}}<{{\tan }^{-1}}x<x"